If you deposit ​$3400 today into an account earning an annual rate of return of 7 ​percent, what would your account be worth in 35 ​years?
Answer #1
A = P*(1+i)^n Where,
= 3400*(1+0.07)^35 A = Future Value = ?
= $ 36,300.38 P = Present Value = $       3,400
i = Rate of return = 7%
n = time = 35